OpenStudy (anonymous):

Solve for x: 2 over 5 plus 3 over 5 x equals quantity x plus 5 over 10

OpenStudy (anonymous):



OpenStudy (anonymous):

whats the LMC?

OpenStudy (anonymous):

LCM*

OpenStudy (anonymous):

i don't know what that is

OpenStudy (anonymous):

I'd multiply every term by 10x the common denominator of the 3 fractions 25×10x+35x×10x=x+510×10x this can be simplified to 4x+6=x(x+5) which is the same as 4x+6=x2+5x or x2+x−6=0 this is a quadratic that can be solved by factoring hope this helps

OpenStudy (anonymous):

lowest common multiple what is the lowest number 5,5,10 has in common

OpenStudy (anonymous):

if x = 3 then (3 - 3) = 0 and 0 * (3 + 2) = 0 same for x = -2 substitute it and you get (-2 -3)(-2 +2) = -5 * 0 = 0

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

@makaylarrrrrr hope i helped

OpenStudy (anonymous):

your welcome

OpenStudy (cj49):

multipy 2/5 by to make the denominater 5x 2x+3/5=x+5/10 cross multiply 10(2x+3)=5x(x+5) 20x+30=5x2+25x -5x2-5x+30=0 -x2-x+6=0 x=-2 or 3